OpenStudy (kmullis6):

hey! is anyone taking algebra 1 on flvs? if so, do you mind helping me? i just started the class a few days ago and i'm terrible at math, especially algebra. thanks!

OpenStudy (anonymous):

I can give it a shot. What problems?

OpenStudy (bakonloverk):

What's the question?

OpenStudy (kmullis6):

Evaluate 6^2 ÷ (3 + 9). 3 4 12 21

OpenStudy (kmullis6):

3?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yeah its 3

OpenStudy (the_fizicx99):

Pemdas, 6^2 = 36 36 / (3+9) 36 / 12 = 3

OpenStudy (kmullis6):

-5 - (3^2 +2) 28 -14 -16 29

OpenStudy (kmullis6):

-16?

OpenStudy (the_fizicx99):

Solve what's int he parenthesis first, solve 3^2, then add 2 to that result. Multiply it by -1 and subtract 5.

OpenStudy (anonymous):

-16

OpenStudy (the_fizicx99):

Yes, -16 is correct.

OpenStudy (anonymous):

-5-(9+2) -5-11=-16

OpenStudy (kmullis6):

Evaluate 12 - 6 + (14 - 8)^2 = ____________. Pretty sure it's 42.

OpenStudy (the_fizicx99):

(14 - 8)^ 2 = 6^2 = 36 12 - 6 = 6 6 + 36 = 42

OpenStudy (anonymous):

42

OpenStudy (kmullis6):

ok thanks you guys. i just wanted to make sure i understood it good enough. (: @tHe_FiZiCx99 @heyitzkelly_marie1484 @Alona

OpenStudy (kmullis6):

Can you help me with 2 more problems? I need help working thru them. I got confused again. /: @heyitzkelly_marie1484 @Alona